μx= 10,    σx =5, μy=8, σy=4,      ρ=0 X and Y come from normal distributions Calculate the...

  1. μx= 10,    σx =5, μy=8, σy=4,      ρ=0

X and Y come from normal distributions

  1. Calculate the probability that X > Y.
  2. Z = 3X + 10   Calculate the probability that Z > 45.
Answer #1

(L-lu) / = ((S-2)/191 8 @ + U: X-Y 3 uy - Anh Học A = 10-8 = 2 53 +57 (P=0) = $25+ 16 = = 541 Za ... Po(x>Y): P(U>0): Plz>-0.